# Eaman Jahani

### v5

@@ 2.2 Decomposition into Model and Prompt Effects
$\Delta Q$ $={}\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{1},s))\;-\;Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Model Effect }(M)\vphantom{\Big{|}}}$ (5)
$+\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{2},s))\;-\;Q(\theta_{2},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Prompting Effect }(P)\vphantom{\Big{|}}}.$
@@
The first term isolates the gain from simply upgrading the model (holding the user’s prompt strategy fixed at the old optimum), while the second term represents any additional gain from prompt adaptation. In other words, even if $\theta$ improves, failing to adjust how prompts are written could leave substantial performance gains on the table.
This decomposition naturally implies a prediction similar to Proposition 1 for each component of $\Delta Q$ that co-evolve with the changes in model capacity and user skill. In particular, when $k$ is small, the following are proved in Appendix E.1.
Proposition 2.
As user skill $s$ improves, the model effect $M$ decreases and the prompting effect $P$ increases.
Intuitively, when costs are low and model capacity rises, higher-skilled users are already near the upper bound of performance, so the direct model effect they receive may be smaller. In contrast, users who can effectively refine prompts might benefit more from the new capacity, so the prompting effect they receive may be larger. Taken together, these theoretical predictions guide our experimental design, where we randomly vary $\theta$ and measure how users adapt.
@@ 3 Experiment Design and Methods
To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted a pre-registered online experiment with 2,059 participants on Prolific between December 12 and December 19, 2023.222Full pre-registration details, including hypotheses and planned analyses, are available in an online repository. Participants were asked to replicate a target image as closely as possible using a generative AI model, with the goal of assessing how both model capability and prompting behavior influence final outcomes. [^3] All procedures were approved by an institutional review board (link removed to preserve author anonymity) and participants provided informed consent. We will release anonymized data and replication code upon publication.
@@ 3.1 Experimental Setting
Each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”). These models differ not only in technical capability but also in whether they apply hidden large language model (LLM)-based modifications to user prompts. In the DALL-E 2 condition, participants interacted with an earlier-generation model that interprets prompts directly without intermediate rewriting. In contrast, the DALL-E 3 API, by default, forwards user prompts to GPT-4 before image generation. This intermediate GPT-4 step rewrites the prompt—typically by adding detail or restructuring language—before passing it to the image model. This behavior is intended to improve image quality but occurs silently and without user visibility. In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered. In the DALL-E 3 (Verbatim) condition, we attempted to suppress GPT-4-based prompt rewriting by prepending a hidden system message instructing the model to leave the user’s prompt unchanged. While some rewriting still occurred, the rate and extent of modifications were substantially reduced. In all conditions, participants were unaware that their prompt might be rewritten or modified before image generation.
In addition to model assignment, each participant was independently assigned one of 15 target images. These images were drawn from three broad categories—business and marketing, graphic design, and architectural photography—to represent common use cases for text-to-image generation. All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography) [^4] All of these images have licenses for free use for commercial and noncommercial purposes.. We built a custom interface resembling ChatGPT, with the assigned target image on the right and a scrollable history of prompts and generated images on the left. Participants were explicitly informed that the model was memoryless: every new prompt was processed independently, carrying no information from previous attempts.
@@
Participants had up to 25 minutes to submit prompts, with a requirement to submit at least 10. They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image was in the top 20% of participants. The median completion time was 22 minutes, implying an average hourly wage of about \$15. After replicating the target image, participants completed a demographic survey covering age, gender, education, occupation, and self-assessed proficiency in creative writing, programming, and generative AI. We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times in a row, or failed to complete the post-task survey, resulting in a final sample of 1,893 participants and 18,560 prompts. [^5] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count is slightly below 10 $\times$ 1,893 because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias. Full details are provided in Appendix C.4.
3.2 Outcome Definition and Stochastic Generation
@@
The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al. 2021). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
@@
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments. [^6] DreamSim (Fu et al. 2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix F. The two measures were highly correlated, and our findings were consistent across both.
@@ 3.3 Replay Analysis for Separating Model and Prompting Effects
A central goal of our experiment was to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts. Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be written as:
We decompose this change into two parts. These are the model effect, the gain from applying the same prompts to a better model,
@@
and the prompting effect, the additional improvement from adapting prompts to take advantage of the more capable model,
@@
To estimate these components empirically, we conducted an additional analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions. To do so, we took the exact prompts that participants submitted during the experiment and re-submitted (or “replayed”) them to both their originally assigned model and the alternative model, generating new images in each case. [^7] Because the replay was conducted several weeks after the original experiment, all prompts were re-evaluated at the same time, using the same infrastructure, regardless of whether they were being run on the original or alternative model. This ensured that any observed differences in output quality could be cleanly attributed to model capacity and prompt adaptation rather than to changes in model behavior over time (i.e., model drift (Chen et al. 2023)). More specifically, prompts written by DALL-E 2 participants—corresponding to $x^{*}(\theta_{1},s)$—were evaluated both on DALL-E 2 and on DALL-E 3, providing an empirical measure of $Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}$ and $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$, respectively. This comparison isolates the model effect: the improvement in output quality when keeping prompts fixed and simply upgrading the model. To estimate the prompting effect, we compared the quality of reused DALL-E 2 prompts on DALL-E 3—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$—to the quality of prompts originally written by DALL-E 3 participants and evaluated on the same model—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}$. This captures the additional improvement from users adapting their prompts to better leverage the capabilities of the more advanced system.
@@ 4 Results
In this section, we present our empirical findings on how increased model capacity affects user performance in replicating target images. In particular, we focus on three main questions: (i) whether access to a more capable model (DALL-E 3) improves performance; (ii) how users modify their prompts in response to the improved model; (iii) and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation. Finally, we compare these empirical results to the predictions of our conceptual framework. Our analysis draws on both the main experimental data and the replay data described above.
4.1 Overall Impact of Model Upgrades
@@
We begin by examining whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2 as implied by Equation 3. Figure 1 summarizes these findings. Panel A presents three representative target images and, for each, three generated images drawn from the full sample of participants across both model conditions. The middle row for each target shows the image whose cosine similarity to the target is closest to the mean similarity across all participants. The rows above (below) show images that are approximately one average treatment effect (ATE) more (less) similar to the target than the mean. This visualization provides qualitative intuition for the magnitude of the effect we estimate: the typical difference in fidelity between participants using DALL-E 2 and those using DALL-E 3 (Verbatim). Panel B shows that, across the 10 required prompt attempts, participants assigned to DALL-E 3 (Verbatim) produce images that are, on average, 0.0164 higher in cosine similarity to the target (95% CI: [0.0104, 0.0224], $p<10^{-5}$). This improvement corresponds to roughly 0.19 standard deviations in performance. The gap persists across all attempts; participants using DALL-E 3 start off producing closer matches and maintain that edge through their 10th prompt.
@@
Figure 1: Overall Performance and Prompting Behavior. (A) For three example target images, the middle row shows participant-generated images closest to the mean similarity across all prompts. The rows above and below show images of approximately one average treatment effect (ATE) more or less similar to the target, illustrating the typical performance difference between model conditions. (B) Top: average CLIP cosine similarity to the target image by attempt, separately for DALL-E 2 and DALL-E 3 participants. Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval. (C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals. Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
@@
Participants’ dynamic prompting behavior also differs substantially between the two models. As shown in panel C of Figure 1, those assigned to DALL-E 3 write prompts that are, on average, 24% longer than those assigned to DALL-E 2, and this gap widens over successive attempts. Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions. Analyses of parts of speech confirm that these extra words likely provide additional descriptive information rather than mere filler: the proportion of nouns and adjectives—the two most descriptively informative parts of speech—is nearly identical across model conditions (48% for DALL-E 3 vs. 49% for DALL-E 2; $p=0.215$). This suggests that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity.
4.2 Replay Analysis and Decomposition of Effects

### v6

@@ 2.2 Decomposition into Model and Prompting Effects
$\Delta Q$ $=\underbrace{\Bigl(Q(\theta_{2},s,x^{*}(\theta_{1},s))-Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr)}_{\text{Model effect }}$ (5)
[…]
@@
The first term isolates the gain from upgrading the model while holding the user’s prompt strategy fixed at its pre-upgrade optimum. The second term captures the additional improvement arising from user-side prompt adaptation. Put differently, even if model capacity improves, one may forgo performance gains if their prompts are not updated accordingly.
2.3 Distributional Impacts of Model Improvements
A related natural question asks whether access to a more capable model levels the playing field, or widens performance gaps between users. In the language of our framework, how does a transition from a model with capacity $\theta_{1}$ to one with higher capacity $\theta_{2}$ affects the distribution of realized output quality. In other words, does access to a more capable model tend to level the playing field, or does it widen performance gaps between users? Under the bounded formulation of the framework analyzed here, $Q(\theta,s,x)=1-e^{-\theta sx}$, improvements in model capacity imply a specific pattern in how performance gains are distributed across users.
Proposition 1 (Equalizing Effect of Model Improvements).
Under the bounded formulation, improvements in model capacity compress the distribution of output quality across users.
This aggregate result can be understood by examining the separate distributional implications of the model effect and the prompting effect.
Proposition 2 (Model Effect: Equalizing Channel).
Holding prompting effort fixed, an increase in model capacity compresses the outcome distribution: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more.
Proposition 3 (Prompting Effect: Expanding Channel).
Allowing users to re-optimize their prompting effort expands the outcome distribution: performance differences across users widen as the upper portion of the distribution shifts further upward.
The formal proofs and derivations for all three propositions are provided in Appendix A.
Taken together, these results imply that model improvements generate two opposing distributional forces. The direct mechanical improvement in capacity—the model effect —is equalizing, while the behavioral response—the prompting effect—is expanding. In aggregate, the equalizing channel dominates, producing the overall compression established in Proposition 1.333In the unbounded but concave logarithmic specification discussed in Appendix A.8, both the model and prompting channels are equalizing rather than opposing. The reversal stems from the additive interaction of capacity and skill, which makes improvements in $\theta$ more uniform across users rather than reinforcing skill differences. Intuitively, as model capacity increases, performance near the top of the distribution is already close to the performance ceiling and experiences smaller marginal gains, while adaptive prompting raises the performance of the upper tail more sharply.
@@ 3 Experiment Design and Methods
To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted two pre-registered online experiments. The first experiment (the image replication experiment, hereinafter) was conducted on a sample of 2,059 participants on Prolific between December 12 and December 19, 2023. In this experiment, participants were asked to replicate a target image as closely as possible using a generative AI model. The second experiment (the logo generation experiment, hereinafter) was conducted on a sample of 2,067 participants on Prolific between June 30 and July 1, 2025. In this experiment, participants were asked to create a logo for a hypothetical organization described in a short textual vignette (e.g., a sign for a business or a crest for a university). In both experiments, the goal was to assess how both model capability and prompting behavior influence final outcomes. [^4] Full pre-registration details for both experiments, including hypotheses and planned analyses, are available in an online repository. All procedures were approved by an institutional review board (protocol information removed to preserve author anonymity) and participants provided informed consent. We will release anonymized data and replication code upon publication.
Together, these two experiments are designed to span distinct but empirically important regimes of generative AI use. The image replication experiment captures a bounded, steerable task in which success can be evaluated against a fixed target and output quality has a natural upper bound. This structure enables precise measurement of performance against an objective benchmark. By contrast, the logo generation experiment captures an open-ended, creative task in which goals are underspecified, quality is inherently comparative rather than absolute, and there is no clear performance ceiling. Despite these differences in task structure, both experiments share a common experimental architecture that allows us to decompose the performance gains into model-driven (model effect) and user-driven (prompting effect) components. Studying both settings allows us to assess whether prompt adaptation operates similarly when users are steering models toward a known objective versus when they are iterating within a creative design space, and to examine how the relative contributions of model capability and user adaptation vary across these regimes.
@@ 3.1 Experimental Setting
In both experiments, each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)” or simply “DALL-E 3”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”). These models differ not only in technical capability but also in whether they apply hidden large language model (LLM)-based modifications to user prompts. In the DALL-E 2 condition, participants interacted with an earlier-generation model that interprets prompts directly without intermediate rewriting. In contrast, the DALL-E 3 API, by default, forwards user prompts to GPT-4 before image generation. This intermediate GPT-4 step rewrites the prompt (typically by adding detail or restructuring language) before passing it to the image model. This behavior is intended to improve image quality, but is not observed by or known to the prompter. In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered.This condition allows us to test whether system-side prompt rewriting serves as a substitute for user-side prompt adaptation, or instead changes users’ ability to steer the model. In the DALL-E 3 (Verbatim) condition, we attempted to suppress GPT-4-based prompt rewriting by prepending a hidden system message instructing the model to leave the user’s prompt unchanged. While some rewriting still occurred, the rate and extent of modifications were substantially reduced. In all conditions, participants were unaware that their prompt might be rewritten or modified before image generation.
Importantly, all generative models in the experiment were memoryless: each prompt was processed independently, with no carryover from prior attempts. Participants were explicitly informed of this property before beginning the task to ensure they understood that previous prompts would not influence subsequent generations. Although this setup departs from typical real-world interfaces that maintain conversational context, it enables us to isolate prompt-level adaptation and to measure users’ iterative learning without confounding from model memory or accumulated dialogue.
@@
We designed a custom experimental interface resembling ChatGPT, which was used for both tasks. The participant’s “objective” (either the target image or logo brief, depending on the task) was displayed on the right side of the screen and a scrollable history of prompts and generated images on the left. For both experiments, participants had up to 25 minutes to submit at least 10 prompts. They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image (according to the relevant outcome metric below) was in the top 20% of participants. The median completion time in the image replication experiment was 23.5 minutes, with an a average hourly wage (including bonuses) of about \$13. The median completion time in the logo generation experiment was 26.1 minutes, with an average hourly wage (including bonuses) of about \$11.50. After they had finished submitting prompts, participants completed a demographic survey covering age, gender, education, occupational skills, and self-assessed proficiency in creative writing, programming, and generative AI. We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times consecutively, [^5] This pre-registered exclusion criterion was intended to filter out low-effort behavior in which participants repeatedly reused the same prompt in an attempt to complete the task as quickly as possible. While repeating a prompt could also reflect legitimate exploration given the stochastic nature of generative AI outputs, we note that this behavior occurred rarely in our data: the criterion affected 15 participants in the image replication experiment and 32 participants in the logo generation experiment (approximately 0.75–1.75% of participants). or failed to complete the post-task survey. This resulted in a final sample of 1,893 participants and 18,560 prompts in the image replication experiment, and a final sample of 1,857 participants and 18,425 prompts in the logo generation experiment. [^6] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count in both experiments is slightly below 10 $\times$ $n_{participants}$ because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias. Full details are provided in Appendix C.4.
3.2 Image Replication Experiment
In the image replication experiment, participants were asked to replicate a “target image” as closely as possible using the generative model provided to them. Each participant was randomly assigned one of 15 target images in addition to their model assignment. These images were drawn from three broad categories—business and marketing, graphic design, and architectural photography—to represent common use cases for text-to-image generation. [^7] The selection of 15 target images was guided by the goal of capturing diversity across common text-to-image use cases (business and marketing, graphic design, and architectural photography), as well as by practical considerations identified through pilot testing. While a larger or systematically pretested image set would be desirable, there is no established standard for sampling text-to-image targets, and for a laboratory experiment of this scope some pragmatic choices were necessary. Our aim was to ensure that tasks were neither trivially easy nor prohibitively difficult, and to align broadly with categories commonly represented on repositories such as PromptBase. Target images were assigned using complete randomization across model–image cells, ensuring that image difficulty was orthogonal to model assignment. In addition, our analysis accounts for variation in difficulty by normalizing performance within each target image. Images were sourced from platforms that permit free research use for commercial and noncommercial purposes (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography). All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography).
3.2.1 Outcome Definition: Expected CLIP Embedding Cosine Similarity
@@
The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (radford2021learning). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
@@
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments. [^8] DreamSim (fu2023dreamsim) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix H. The two measures were highly correlated, and our findings were consistent across both.
3.3 Logo Generation Experiment
In the second experiment, each participant was asked to generate an original logo based on a short set of instructions or “briefs.” Each participant was randomly assigned one of 15 briefs in addition to their model assignment. One example is:
“The owners of Green Pine Brewery are seeking a logo for their family-owned craft brewery located in Vermont.
  • Required visual elements: pine tree, hops, barley
  • Context for use: beer labels, merchandise, digital marketing
Please create the best possible design based on the client’s description and the listed requirements.”
All briefs followed this exact format: the name of the organization, a single sentence of background information, 2–3 required visual elements, and the anticipated context of use.
The briefs span a range of topics, from a community park sign to a soccer team crest to a university letterhead. We provide the full list of briefs in Appendix B. They were written with three goals in mind: (i) capture a wide variety of realistic use cases, (ii) include a small number of concrete required elements that provide shared reference points for evaluators in an otherwise subjective task, and (iii) allow for open-ended creative possibilities. This third feature contrasts with the image replication experiment, where the best possible performance is well defined (perfectly replicate the target image).
3.3.1 Outcome Definition: Bradley-Terry Strength
Because each logo-generation task was open-ended and lacked an objective reference image, we could not assess output quality using embedding-based similarity metrics (e.g., cosine similarity of CLIP representations). Instead, we employed the Bradley–Terry (BT) model (bradley1952rank), a standard approach for inferring latent quality scores when only relative judgments are available (zermelo1929berechnung, ford1957solution). At a high level, the BT model is a probabilistic model that estimates a strength parameter (which we refer to hereafter as “BT $\beta$”) for each item based on the outcomes of many pairwise comparisons.
[…]
@@ 3.4 Replay Analysis for Separating Model and Prompting Effects
A central goal of our experiment is to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts. Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be decomposed into two parts: the model effect, i.e., the gain from applying the same prompts to a better model,
@@
and the prompting effect, i.e., the additional improvement from adapting prompts to take advantage of the more capable model,
@@
To estimate the model and prompting components empirically, we conducted a replay-based analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions. In both experiments, we re-evaluated prompts under the alternative model—replaying DALL-E 2 prompts on DALL-E 3 and DALL-E 3 prompts on DALL-E 2—thereby holding prompt content fixed while varying model capacity. These cross-model evaluations provide empirical analogues of $Q(\theta_{2},s,x^{*}(\theta_{1},s))$ and $Q(\theta_{1},s,x^{*}(\theta_{2},s))$, respectively, and together with the original prompt–model pairings, allow us to implement the decomposition into model and prompting effects shown in Equation 5.
In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (chen2023gptchanging). In the logo generation experiment, replay occurred shortly after data collection, so this additional step was unnecessary.
@@ 4 Results
In this section, we present our empirical findings on how increases in model capacity affect user performance across the two tasks. For each task, we first examine whether access to a more capable model improves average output quality, how users adapt their prompts in response, and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation using the replay analyses. Next, we study how these gains reshape the distribution of outcomes across users and interpret these distributional impacts through the lens of our conceptual framework. Finally, we assess the impact of automated prompt rewriting using the DALL-E 3 Revised condition.
4.1 Image Replication Task
We begin with the image replication task. As discussed above, this task represents a bounded, steerable application of generative AI, in which users aim to match a known target and output quality has a natural upper bound.
4.1.1 Overall Impact of Model Upgrades
@@
We first examine whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2, as implied by Equation 3. Figure 1 summarizes these findings. The top row of Panel A presents three of the target images provided to participants. Below each target, three generated images are shown, drawn from the full sample of participants across both model conditions. The middle row for each target shows the image whose cosine similarity to the target is closest to the mean similarity across all participants. The rows above (below) show images that are approximately one average treatment effect (ATE) more (less) similar to the target than the mean. This visualization provides qualitative intuition for the magnitude of the effect we estimate: the typical difference in fidelity between participants using DALL-E 2 and those using DALL-E 3 (Verbatim). Panel B shows that, across the 10 required prompt attempts, participants assigned to DALL-E 3 (Verbatim) produce images that are, on average, 0.0164 higher in cosine similarity to the target (95% CI: [0.0104, 0.0224], $p<10^{-5}$). This improvement corresponds to roughly 0.19 standard deviations in performance. The gap persists across all attempts; participants using DALL-E 3 start off producing closer matches and maintain that edge through their 10th prompt.
Figure 1: Overall Performance and Prompting Behavior for the Image Replication Task
@@
(A) For three example target images, the middle row shows participant-generated images closest to the mean similarity across all prompts. The rows above and below show images of approximately one average treatment effect (ATE) more or less similar to the target, illustrating the typical performance difference between model conditions. (B) Top: average CLIP cosine similarity to the target image by attempt, separately for DALL-E 2 and DALL-E 3 participants. Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval. (C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals. Color shading indicates the average textual embedding similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
@@
Participants’ dynamic prompting behavior also differs substantially between the two models. As shown in panel C of Figure 1, those assigned to DALL-E 3 write prompts that are, on average, 24% longer than those assigned to DALL-E 2, and this gap widens over successive attempts. Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale—see Appendix C.3 for details), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions. Analyses of parts of speech confirm that these extra words likely provide additional descriptive information rather than mere filler: the proportion of nouns and adjectives—the two most descriptively informative parts of speech—is nearly identical across model conditions (48% for DALL-E 3 vs. 49% for DALL-E 2; $p=0.215$), suggesting that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity. Together, these patterns illustrate prompt adaptation in practice: users progressively supply more information-dense prompts as they naturally learn the model’s affordances—despite no explicit instructions to do so.
4.1.2 Replay Analysis and Decomposition of Effects